OpenStudy (anonymous):

Find the real numbers x and y that make the equation true. 5 + yi = x + 3i

OpenStudy (anonymous):

\(a+bi=c+di\iff a=c, b=d\) so in your case \(5=x\) and \(y=3\)

OpenStudy (anonymous):

Just Compare real & imaginary parts seperately x=5 & y=3 In general a+ib=c+id when a=c && b=d